OpenStudy (anonymous):

How do you make those schnazzy fractions that actually look like fractions? I'd like a fraction to appear as an a over b instead of a/b...

OpenStudy (anonymous):

Same question

myininaya (myininaya):

frac{something}{something}

OpenStudy (saifoo.khan):

good ques.

OpenStudy (saifoo.khan):

frac{saif}{saif}

OpenStudy (anonymous):

\[\frac{awesome}{sauce}\]

OpenStudy (anonymous):

Great! Thank you!

OpenStudy (saifoo.khan):

it didnt worked! :O

OpenStudy (saifoo.khan):

frac{saif}/{saif}

myininaya (myininaya):

type it into equation thingy below lol

OpenStudy (saifoo.khan):

it aint working!

OpenStudy (saifoo.khan):

\frac{saif}{saif}

OpenStudy (saifoo.khan):

\frac{saif}{saif}\

OpenStudy (anonymous):

Hit the "Equation" button first

myininaya (myininaya):

khan do you see equation button

OpenStudy (saifoo.khan):

\[\frac{saif}{saif}\]

myininaya (myininaya):

gj :)

OpenStudy (saifoo.khan):

\[\frac{yay}{christey_rocks!}\]

myininaya (myininaya):

christy*

OpenStudy (anonymous):

If you can get the desired format for your output, you might try creating a document file and attakh it by clicking the blue Attach Kile button below.

OpenStudy (saifoo.khan):

\[\frac{Correction}{Christy}\]

OpenStudy (anonymous):

Thanks all!

myininaya (myininaya):

\[\int\limits_{christy}^{jabber}khan dkhan=(khan)^2|_{christy}^{jobber}\]

myininaya (myininaya):

oops divided by 2 lol

OpenStudy (saifoo.khan):

\[\frac{lol}{good one}\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\frac{3}{4}\[\frac{3}{4} \]